OpenStudy (jotopia34):

Help, quiz tomorrow andIdon't know how to do the Integral of tanx*ln(cosx) dx

OpenStudy (jotopia34):

\[\int\limits_{0}^{\pi/3}tanx*\ln(cosx)dx\]

OpenStudy (anonymous):

\[\int_0^{\pi/3}\tan x \;\ln(\cos x) \;dx\\ \int_0^{\pi/3}\frac{\sin x}{\cos x} \;\ln(\cos x) \;dx\] Have you tried a substitution?

OpenStudy (jotopia34):

I did try a u sub with u=cos x

OpenStudy (anonymous):

that should work... ^^^

OpenStudy (jotopia34):

I got du= -sinx dx cancelled the sin x , but then couldntdo the 1/u ln(u) integral :(

OpenStudy (jotopia34):

help!!!!!!!

OpenStudy (anonymous):

try u =ln(cosx)

OpenStudy (anonymous):

So, using the substitution \[u=\cos x,\\ du=-\sin x \;dx\\ -du=\sin x\; dx,\] the integral becomes \[-\int_1^{\frac{1}{2}}\frac{\ln u}{u}du\\ \int_{\frac{1}{2}}^{1}\frac{\ln u}{u}du\] Try another substitution. @joemath314159's sub is quicker, but I've a preference for step-by-step subs.

OpenStudy (jotopia34):

I am gonna try u=ln(cosx)

OpenStudy (jotopia34):

I got it, thank you joemath and giggles. I always feel like you cant do an f(g(x)) u substittution